OpenStudy (anonymous):

Find a number k satisfying the given condition. x + 4 is a factor of x3 - kx2 + 2x + 4k

OpenStudy (anonymous):

substitute x=-4 in the equation \[=>(-4)^3+8k-8+4k=0\]\[=>-64+12k-8=0\]\[=>k=\frac{72}{12}=6\]

OpenStudy (anonymous):